OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.2 megabytes (MB). The size of the high-quality version is 4.7 MB. Yesterday, there were 1400 downloads of the song, for a total download size of 5105 MB. How many downloads of the standard version were there?

OpenStudy (shane_b):

Set up the equations:\[2.2x+4.7y=5105\]\[x+y=1400\]Can you solve it from here?

OpenStudy (anonymous):

no sorry

OpenStudy (shane_b):

Ok...it's two equations with 2 unknowns. Basically take one variable and solve it for the other. The second equation is simpler...so the first step is to solve it for x or y (it doesn't matter). Can you do that part?

OpenStudy (shane_b):

(just rearrange it so that x or y = ...)

OpenStudy (anonymous):

no dont understand

OpenStudy (shane_b):

Ok...let's solve x in terms of y:\[x+y=1400\]\[x=1400-y\] With me so far?

OpenStudy (anonymous):

ya

OpenStudy (shane_b):

Great. Now just plug that in anywhere you see x in the first equation. The first equation was \[2.2x+4.7y=5105\]Now it becomes: \[2.2(1400-y)+4.7y=5105\]

OpenStudy (shane_b):

Now it's possible to solve for y. Once you have y, you can solve for x in the original equations.

OpenStudy (anonymous):

im looking for the answer]

OpenStudy (anonymous):

for y i got 810

OpenStudy (anonymous):

how do you solve for x

OpenStudy (shane_b):

810 is correct...you're almost done. Remember that one of the original equations was x+y = 1400. Well, now you have y...just solve for x.

OpenStudy (shane_b):

\[x+810=1400\]x=?

OpenStudy (anonymous):

x=590

OpenStudy (shane_b):

Perfect...you're done. Good luck :)

OpenStudy (anonymous):

so is 590 the answer? thank you for your help

OpenStudy (shane_b):

x=590, y = 810

OpenStudy (shane_b):

yw